OpenStudy (anonymous):

b+9 _____ =-3 what is b? step-by-step. 12

OpenStudy (owlcoffee):

we have: \[\frac{ b+9 }{ 12 }=-3\] Now, algebraically that means a "proportion", so we'll cross-multiply: \[b+9=-3(12)\] The -3 has an invisible "1" below it, so multiplying a number by one gives me the same number. Let's operate that multiplication: \[b+9=-36\] Now, let's substract 9 to each side to get: \[b+9-9=-36-9\] the 9s will simply so we'll end up with: \[b=-36-9\] We know that if we substract a number from a negative number, is the same as summing the numbers in a negative form: \[b=-45\] And we've found the value of b.